40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15245     Entry time: Tue Mar 3 19:11:48 2020
Author: gautam 
Type: Update 
Category: LSC 
Subject: Some locking prep 
  • Re-aligned and locked the arm cavities for IR and green.
  • Re-set trans normalization because after the c1iscex and c1iscey reboots, these didn't come back to the old values.
ELOG V3.1.3-